ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60F05 60F15 Rempala, Grzegorz A., University of Louisville, KY USA Gupta, A. K. Bowling Green State University, OH USA Random permanents of increasing degree Let Ξ_{m,n} = [ξ_{ij}] be a random m × n (m ≤ n) matrix of independent entries bounded with probability one and satisfying sup_{i,j} ∑_{k=1}^{m} |Eξ_{kj} −Eξ_{ki}|/m → 0 as n, m → ∞. In this work we derive the limiting laws for Per Ξ_{m,n} as m, n → ∞ and m/n → λ > 0. Some examples of applications of our results in statistical inference involving multiparameter multinomial distribution are given.